OpenStudy (anonymous):

Which choice shows the next step needed to simplify 6(12) - 6(5) using the Distributive Property? A. 6(12 - 5) B. 12(6 - 5) C. 12(12 - 5) D. 5(12 - 6)

OpenStudy (nubeer):

take 6 common from ur given equation ... it will give ur answer

OpenStudy (jiteshmeghwal9):

by taking common 6

OpenStudy (ash2326):

Distributive Property \[A(B-C)=A\times B-A\times C\] here we are using the opposite of it to simplify. Could you check which options uses that? @jamesballsfacenow